Suppose when disposable income (D) is $1000/month, consumption (C) is $400/month. When disposable income is $2500/month, consumption increases to $700/month. Which of the following equation represents the relationship between consumption and disposable income in this example?
Answer #1

Marginal Propensity to consume is a measurement indicator which calculates that an increase in personal consumption expenditure is accompanied by an increase in income.

Marginal Propensity to consume is represented as :

MPC = Δc/ΔΥ

MPC = Change in Consumption / Change in Income

MPC = 700 - 400 / 2500 - 1000

= 300 / 1500

= 1/5

or 0.2

So Marginal Propensity to Consume = 0.2